OpenStudy (anonymous):

Can anyone help me with about 25Qs i have the flu and were doing a section i dont know how to and im already behind from the flu plz!!!

OpenStudy (anonymous):

Marshall is writing an equivalent equation to solve for x. What should be his next step? 5x+4y=8 5x+4y-4y=8-4y 5x=8-4y

OpenStudy (anonymous):

and the options are Add 5x to both sides of the equation. Subtract 5x from both sides of the equation. Multiply both sides of the equation by 5. Divide both sides of the equation by 5.

OpenStudy (anonymous):

A regular hexagon and a rectangle have the same perimeter, P. A side of the hexagon is 4 less than the length, l, of the rectangle. The width of the rectangle, w, is 2 less than the length of the rectangle. What is the perimeter of the hexagon? es025-1.jpg

OpenStudy (anonymous):

last one plz

OpenStudy (anonymous):

what about after i close it you can still help?

OpenStudy (anonymous):

its closed now can you answer that one i just asked

OpenStudy (anonymous):

12/5

OpenStudy (anonymous):

hey, that's all? no more?

OpenStudy (anonymous):

at like 1;15 ill need more help